Documentation for TimeGuardian

Version 1.0, 4.7.94 written by Markus Matern

translated by Markus Matern and Gerri Körner

This is the documentation for the package TimeGuardian Version 1.0, 4.7.94. The descriptions refers to the following programs:

• TGCron Version 1.42, 4.7.1994

• TimeGuardian Version 1.40, 4.7.94.

Gerri Körner

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994 by Gerri Körner

1 Introduction

Welcome to the documentation of the program package TimeGuardian.

This package may be freely distributed, as long as no charges more than reasonable copying and handling fees are collected. For any other type of distribution, you must have the written permission of the author. This package is shareware, when you want to use the programs of this package permanently, you have to register, see Section 1.5 [How can I register for TimeGuardian?], page 3 for further information.

This program may be included in software collections, providing that the previous conditions are respected.

This program package is provided without warranty of any kind. In no event will the author be liable for direct, indirect, or incidental damages resulting from any defect of the program or in its documentation. The users are hereby warned of the possibility of such damage occurences.

1.1 What is TimeGuardian?

TimeGuardian is a program package with which you can easily start events, that have to be done at certain times or dates on your computer. You also can write log files to disk, where you can see when your computer was switched on.

1.2 The features of TimeGuardian

• easy to configure

• Alarm

• daily actions

• yearly actions

• events in time intervals

• localized gui

  • Installer

• MUI-GUI

• AREXX

• online help

1.3 What can I do with TimeGuardian?

The package has two main functions, you can use seperately from each other.

The first function it provides, is a function to save on disk every time your computer was turned on and off. You get a complete log file when anyone turned on your computer. When you want, you are able to log different projects. That means, that you see in the log files, when you wirked on a certain project, e.g. private work and business work. You can analyse these log files easily with the graphical user interface, so you get a quick overview when you spend time at your computer or at certain projects.

The second function, that is provided by this package is a function to start actions (commands or scripts) that you want to be executed regularly or at specified dates or hours. Moreover you can start programs or scripts once a day or once a year.

1.4 What do I need, to be able to use TimeGuardian?

When you want to use all features of the package TimeGuardian, you need some other public domain or shareware packages, which can’t be included in this package, because of different reasons.

• OS version >= 2.0 You need a Workbench or Kickstart version greater or equal to 2.0 by all means. When you have not installed this two versions you can’t install this package. When you want to know, which versions you have installed, enter in a ‘Shell’: ‘version’. The result should be like follows:

‘Kickstart version 37.175, Workbench version 37.XX’ or greater numbers! You need a

Workbench versioon >= 2.1 when you want to use the localized graphical user interface!

• MUI >= 2.0 You need the program package MUI to be able to use the graphical user interface of the program TimeGuardian. You can use an unregistered version of MUI when you are testing the program package. When you want to use TimeGuardian regularely, then please register for both packages MUI and TimeGuardian. For further information on how to do this, see Section 1.5 [How can I register for TimeGuardian?], page 3 and Section 3.1 [Where can I obtain MUI?], page 30.

• Installer

You need this program to install the package easily. Many other program packages distribute this program. Where you can obtain the program Installer, see Section 3.2 [Where can I obtain Installer?], page 30.

  • Amiga-Guide

You need this program, that you can use a feature of the graphical user interface of TimeGuardian, the online help. When you want to get information, on how to obtain this package, see Section 3.3 [Where can I obtainAmiga-Guide?], page 31.

1.5 How can I register for TimeGuardian?

That you are able to use the background program TGCron as long as you want, on a certain day, you need a certain file, you only get, when you register. This file containd your name and is generated for you personally. When this file is missing, you can use the program only for about one and a half hour a day.

To obtain your personal key file, please send 20.- DM or 15.- US-$ in notes and the registration form to one of the below adresses (please note, that only assigments from german users to a german bank are accepted, because the fees for assignments from foreign banks are as high as the registration fee). When you send the money, be sure that you can’t see the notes inside! Put the notes inside the registration form.

Gerri Körner Kulturenweg 6 D-87642 Buching

Markus Matern Max-Wieland-Strasse 13 D-89134 Blaustein

You get your personal key file and the newest version of the program package on disk.

You pay for the actual version of the program package TimeGuardian and have neither the pretension to get newer, nor to get enhanced program versions. You don’t have to register again for possible new versions of this package. The personal key files will stay valid with new versions. Possible enhancements of the package will depend on the number of registrations.

We both can’t be reached via telephone the next time, so when you have questions, please send an international reply coupon to one of the above adresses. The better way is to send an E-Mail to (other forms of questions unfortunately can’t be answered.):

               ‘Gerri.Koerner@student.uni-ulm.de’ or
                ‘Markus.Matern@student.uni-ulm.de’.

2 The programs of the package TimeGuardian

The package TimeGuardian consists in general of two programs, you need both to run properly. They will be described later on. You need the preferences program TimeGuardian only, when you want to change the configuration of the cron^{1}, which is running in the background. Or you have to start the first program, when you want to analyze the files, the package writes on your disk. The second program is the cron TGCron. You should start it every time your computer is booting, when you want to use the package in a sensible way. See Chapter 4 [How can I install TimeGuardian?], page 32, for further information.

2.1 The program TimeGuardian

The preferences program TimeGuardian is designed, to help the user to modify the configuration of the background program easily. You have a graphical user interface to do this. When you want to change the configuration, just start TimeGuardian.

• From Workbench: Doubleclick on the corresponding icon. When you followed the proposals of the installation, you find the icon in your drawer ‘Prefs’ of your bootpartition.

• From CLI: Just enter ‘<path>/TimeGuardian’. Instead of <path> you have to enter the name of the drawer in which you have installed the package TimeGuardian. When you followed the proposals of the installation this is ‘Sys:Tools’.

Now follows a description of the different elements of the graphical user interface and their functions. Gadgets are elements, that start an action immediately. They can be activated by clicking on them with the mouse or by pressing the underlined character of their name. You always have to press a single key, even when a capital letter is underlined. Pages are gadgets that activate more and other gadgets. Depending on the preferences of MUI you can see all the page-gadgets together or one after the other by activating the visible page-gadget.

1 A cron is a programm, that starts user-specified programs or scripts at certain user-specified moments.

2.1.1 The menus of the program TimeGuardian

Some program-functions can be activated with gadgets and menus. The menus are shown further on in this documentation. The functions, that are available via gadget and menu will be described at the corresponding gadgets. The menu structure of the preferences program looks like follows:

• Project − Open − Save as − ————— − User − Remove Cron − Iconify − ————— − Quit

• Edit − Reset to Defaults − Last Saved − Restore

• Settings − Create Icons?

• Pages − About − Path Setup − Cron Setup − Global Statistics − Daily Statistics

2.1.1.1 The menu ‘Project⇒Open’

This menu opens a file-requester, with which you can select a configuration file for loading. You can choose a file saved before, with any name. If you don’t select a TimeGuardian- config-file an errormessage will appear. Is the file loaded succesfully, the gadgets and string gadgets will show the new configuration. Now you are able to modify, save, use, these new settings in the custom manner.

2.1.1.2 The menu ‘Project⇒Saveas’

When you activate this menu you have the possibility to save your current settings under any name you choose. The configuration will not be saved in a way, that it is used immediataly. You have to select the gadget ‘Save’ or ‘Use’ when you want to do this. This menu is provided to be able to change quickly between often used configuratons (e.g. one configuration for private work and one for business work). Each configuration has its own name, which easily can be associated with the corresponding project it is desired for. When you have activated the menu ‘Settings⇒Create Icons?’ there will be generated an icon for every configuration you save. When you want to use one of your saved configurations you have to doubleclick on the corresponding icon. This configuration is only used as long as you don’t boot. When you want to use this configuration permanently, when you doubleclick the icon, then you have to change the ‘Tool Types:’ of the icon to ‘ACTION=SAVE’. To do this, activate the icon (don’t doubleclick it!) and select the menu ‘Icons⇒Information..’ on the Workbench. Then change the line ‘ACTION=USE’ to ‘ACTION=SAVE’ and press ‘Save’.

2.1.1.3 The menu ‘Project⇒User’

When you activate this menu you are shown an information window, in which you can read the message

This TimeGuardian is registered to:

Firstname Lastname Street Post code Location Country

when you are a registered user, or the message

This is a unregistered version of

TimeGuardian

you have to register after an evaluation time of 14 days!!

when you haven’t registered yet. When you want to know how to register for TimeGuardian, see Section 1.5 [How can I register?], page 3.

2.1.1.4 The menu ‘Project⇒Iconify’

When you activate this menu, the program closes the window of the graphical user interface and appears as an icon on the Workbench. When you want to reactivate the window, you have to doubleclick the icon. This menu has the similar effect like the third gadget from the right, on the top window border.

2.1.1.5 The menu ‘Edit⇒LastSaved’

When you activate this menu, the configuration file that you saved with the gadget ‘Save’ will be loaded. When the file is loaded succesfully you can edit, use, save it in the common manner.

2.1.1.6 The menu ‘Settings⇒CreateIcons?’

When this menu has a checkmark, each time you save a configuration with the menu ‘Project⇒Save As’, an icon for the configuration will be created.

2.1.2 The page ‘About’

This page shows the copyrights, the shareware reference and the reference to MUI. It is shown when you start TimeGuardian and when you make one of the folowing actions on the graphical user interface:

• Activate the gadget ‘About’

• Select the menu ‘Pages⇒About’

• Press the keys AMIGA 1

• Press the keys RALT SPACE, when you are on the page ‘Daily Statistics’

• Press the keys LALT SPACE, when you are on the page ‘Path Setup’

2.1.3 The page ‘PathSetup’

This page is shown, when you make one of the folowing actions on the graphical user interface: • Activate the gadget ‘Path Setup’ • Select the menu ‘Pages⇒Path Setup’

• Press the keys AMIGA 2

• Press the keys RALT SPACE, when you are on the page ‘About’

• Press the keys LALT SPACE, when you are on the page ‘Cron Setup’

On this page you can select all the files and paths TGCron accesses. You see eight different string gadgets, in which you can enter the values on different ways:

• Click with the mouse inside the frame of the desired string gadget.

• Press the underlined character of the desired string gadgt on the keyboard.

• The string gadgets are concated, press the key that you selected in the MUI preferences program for concatenation (TAB, when you didn’t change this option). Press this key so many times, that the cursor appears in the desired string gadget.

Enter the desired Command or Path on your keyboard now. In the first and the last four string gadgets the entered values will be checked. When they don’t exist they will be rejected.

• Press the gadget in the right of the string gadget. You then will be able to select the path or command/ script with a file requester.

When you start TimeGuardian the first time, you will find some values filled in, in some string gadgets. Otherwise you will see the last saved configuration.

Besides you see four gadgets, with which you can make different changes to the configuration. For further information see Section 2.1.5 [The gadgets ‘Save’/ ‘Use’], page 20 and Section 2.1.6 [The Gadgets ‘Default’/ ‘Restore’], page 20.

2.1.3.1 The string gadget ‘LogPath’

In this string gadget you have to enter the drawer or partition, where the cron program shall save, when the computer has booted and how long and when it was switched on. The names of these three files you can enter in the following string gadgets.

When you want to change the path, activate the string gadget, like described above, and enter the desired path. The path must already exist, otherwise it is rejected. When you activate the string gadget with the gadget in the right of it, a file requester pops up, where you easily can select the path.

When you press the gadget ‘Default’ the path ‘Sys:Tools/’ appears.

2.1.3.2 The string gadget ‘Starts’

Here you have to enter the name of the file, in which TGCron shall save the time, when the computer has had a reset. The file will be created in the drawer described above, when it doesn’t exists. The suffix is automatically created by the program and is a period, followed by the year (in the year 1994: ‘.94’).

When you select the gadget ‘Default’ the file name ‘Logfile’ appears.

The values in the file will only show your real boot times, when you start the program TGCron every time your computer is booting. How you can reach this, see Chapter 4 [How can I install

TimeGuardian?], page 32.

When you start TGCron manually, you can each time you work on a certain project, start the program. You then have a protocol for this specific project. You can reach this in a second way. You start the program TGCron always and change the configuration for different projects. You define a configuration for a certain project and save it with the menu ‘Project⇒Save As’. When you start working on the project you use the menu ‘Project⇒Open’ to load it. For further information, see Section 2.1.1.1 [The menu ‘Project⇒Open’], page 5 and Section 2.1.1.2 [The menu ‘Project⇒Save As’], page 6.

The file is saved in ASCII format and can be loaded with any editor or text viewer. A sample file named ‘Logfile.94’, could look like this:

     Wed   04-May-94 15:26
     Thu   05-May-94  8:24
     Thu   05-May-94 14:26
     Thu   05-May-94 15:22
     Thu   05-May-94 16:01

The date format can be changed, see Section 2.1.4.3 [The gadget ‘Date Format’], page 18 for furher information.

2.1.3.3 The string gadget ‘Runtime’

Here you have to enter the file, in which TGCron shall save how long the computer was switched on at which dates. When the file doesn’t exist, it will be created in the drawer or partition, you selected in the first string gadget on this page. The suffix is automatically created by the program and is a period, followed by the year (in the year 1994: ‘.94’).

When you select the gadget ‘Default’ the file name ‘Uptime’ appears.

The entry for a day will be written the following day, you switch your computer on. That means, when you look at this file the actual day isn’t shown.

The file is saved in ASCII format and can be loaded with every editor or text viewer. A sample file named ‘UpTime.94’, could look like this:

     Wed   04-May-94 Runtime:  2 h  47 min.
     Thu   05-May-94 Runtime:  2 h  43 min.
     Mon   09-May-94 Runtime:  3 h  11 min.

The date format can be changed, see Section 2.1.4.3 [The gadget ‘Date Format’], page 18 for furher information.

2.1.3.4 The string gadget ‘OnOff’

Here you have to enter the file, in which TGCron shall save the times when the computer was switched on and off. When the file doesn’t exist, it will be created in the drawer or partition, you selected in the first string gadget on this page. The suffix is automatically created by the program and is a period, followed by the year (in the year 1994: ‘.94’).

When you select the gadget ‘Default’ the file name ‘OnOff’ appears.

The entry for the current switch on will be written to the file, the next time, the computer is switched on. When you look at the file, the last time you switched the computer on, may be saved

like follows: System switched on: Th 05-May-94 14:26.

The file is saved in ASCII format and can be loaded with any editor or text viewer. A sample file named ‘UpTime.94’, could look like this:

     System switched on:   Wed   04-May-94 17:38
     Resets:  0
     System switched off:  Wed   04-May-94 19:26
     System switched on:   Thu   05-May-94  8:24
     Resets:  0
     System switched off:  Thu   05-May-94  9:25
     System switched on:   Thu   05-May-94 14:26
     Resets:  2
     System switched off:  Thu   05-May-94 16:11

The date format can be changed, see Section 2.1.4.3 [The gadget ‘Date Format’], page 18 for furher information.

The decision, whether the computer is switched on or had a reset, is made depending on the settings of the slider ‘Startuptime’, see Section 2.1.4.1 [The slider ‘Startuptime’], page 18, for further information.

2.1.3.5 The string gadget ‘Onceaday’

Here you can select a file that is executed once a day. The file is executed only once a day independent of the time, the computer is switched on. When the computer works at midnight, it will be executed short after midnight. This entry is not fit for polling at your mailbox once a day, because the time when this happens is not determined. It is intended for commands which you don’t want to execute from your ‘Startup-Sequence’ every time your computer is booting. You can make a backup once a day of important files or update your database for a fast file search.

When you press the gadget in the right of the string gadget, a file requester pops up, where you easily can select the desired file. You can choose an executable or a script in ASCII format. When you want a script to be executed, the ‘S-flag’ has to be set. For example the script ‘Once_a_day’ in the drawer ‘S:’. Enter ‘protect S:Once_a_day s add’ in a CLI to set this flag. You have to start commands inside scripts with their full path! You don’t need the path, when you use intern commands or commands in the ‘C:’-directory.

A sample script may look like this:

     ;Skript once a day
     dir all sys: > T:.directorytree ;for fast file search
     copy T:.directorytree sys:
     delete T:.directorytree

The comands will be automatically started in the background. You don’t have to enter anything like ‘run < nil: > nil: Command’. Bevor adding a command line to a script, you want to execute, try the command in a ‘Shell’ so that you can be sure it works the intended way. When this is not successfull you can start the command from the script and redirect the output to a file: ‘Command > Sys:.logfile’. You can read the file ‘Sys:.logfile’ with a text viewer or an editor to look what went wrong (you can look at this file with ‘type Sys:.logfile’ in a ‘Shell’, too).

2.1.3.6 The string gadget ‘Onceayear’

Here you can select a file, that is executed once a year. The file is executed only once a year, independent of the time, when the computer is switched on the first time in a year. When the computer works at the turn of the year, it will be executed short after midnight. You could use this entry to copy the files, the cron program produced, to another directory. In the drawer, you selected in the string gadget ‘Log Path’, you will only find the actual files.

When you press the gadget in the right of the string gadget a file requester pops up, where you easily can select the desired file. You can choose an executable or a script in ASCII format. When you want a script to be executed, the ‘S-flag’ has to be set. For example the script ‘Once_a_year’ in the drawer ‘S:’. Enter ‘protect S:Once_a_year s add’ in a CLI to set this flag. You have to start commands inside scripts with their full path! You don’t need the path, when you use intern commands or commands in the ‘C:’-directory.

An example script, that copies the files from the drawer ‘Sys:Tools’ to a new drawer ‘Sys:Tools/Store’ could look like this (the drawer ‘Store’ will be created, when the drawer ‘Sys:Tools’ already exists):

     ;Script once a year

copy Sys:Tools/LogFile.?? Sys:Tools/Store quiet
delete Sys:Tools/LogFile.?? quiet
copy Sys:Tools/UpTime.?? Sys:Tools/Store quiet
delete Sys:Tools/UpTime.?? quiet
copy Sys:Tools/OnOff.?? Sys:Tools/Store quiet
delete Sys:Tools/OnOff.?? quiet

The used names and paths, are the ones, that you get, when you press the gadget ‘Default’. When you have chosen other names and want to use this script, then you have to modifiy the script to suit your needs.

2.1.3.7 The string gadget ‘Timerfixed’

Here you can select a CronTab^{2} file in which is saved which programs have to be started at certain moments. This file has to be an ASCII file with a certain format, that will be described below. It is possible to use samples in this file or certain days, months, etc.

When you press the gadget in the right of the string gadget, a file requester pops up, where you easily can select the desired file. You can only access this string gadget, when you have selected ‘fixed’ or ‘fixed & interval’ at the gadget ‘Type’ on the page ‘Cron Setup’. Otherwise the string gadget is ghosted.

An example file could look like follows:

; TimeGuardian CronTable
;
;55 23 * * * RequestChoice TimerRequest "It is  23:55 " "What else ?!"
;
59 19 * * 1-5 s:News
;
0,15,30,45 * * * * "newshell con:82/175/550/90/UpTime from batch:saytime"
;

2 The format is the Un*x format When the 1st character of a line is ‘;’, the line will be ignored completely. You can use this to add comments to the file or to stop executing a program for some time. This was done in the third line. The command ‘RequestChoice’ won’t be executed.

Shall a file be executed at a certain moment, the line in the file must look like the folowing

sample: ‘<Minute> <Hour> <Day> <Month> <Day_of_week> <Path/<Command|Script>>’. The de-

cision, whether a command is executed or not, is made by a AND- combination of the five time parameters. That means, that the command is only executed, when all five parameters are true.

Now follows an exact description for the different samples, set in acute parenthesis. Some characters have a certain meaning:

− ‘<>’:

The parameter in acute parenthesis has to be specified. − ‘|’:

The vertical bar seperates alternative parameters, which can be used either or. − ‘[]’:

The parameter in angular parenthesis can be used when you want. − ‘…’:

Three periods mean, that the parameter before can be used multiple times.

• ‘<Minute>’

The complete sample looks like this:

‘<*|Min[[,Min]|[,Min-Min]. . .]|Min-Min[[,Min]|[,Min-Min]. . .]>’

The value for ‘Min’ has to be choosen in the range from 0 to 59. It follows a description of all possibilities:

1. ‘*’:

When you use an asterisk for this parameter, th command is executed every minute.

2. ‘Min’:

You can choose a moment, at which the command is executed.

 3. ‘Min-Min’:

You can specify a range, whilst which the command is executed. The first value for ‘Min’ has to be smaller or equal than the second value for ‘Min’.

 4. ‘<Min>,<Min-Min>,. . .,<Min>’:

You can specify multiple moments and ranges seperated by colons. The command is executed at every specified moment and whilst the ranges.

Example: ‘0,5,10,15-20,30,45-50 * * * * Command’

That would start the command every full hour, five, ten and 30 minutes after the full hour and in the range from 15 to 20 and from 45 to 50 minutes every minute after the full hour. The command is executed every day.

• ‘<Hour>’

The complete sample looks like this:

‘<*|Hour[[,Hour]|[,Hour-Hour]. . .]|Hour-Hour[[,Hour]|[,Hour-Hour]. . .]>’

The value for ‘Hour’ has to be choosen in the range from 0 to 23. For a more precise description of the parameter ‘Hour’ look at the parameter ‘Minute’ above. Both parameters have the same sample.

Example: ‘* 0,3,6,9,12,15,18,21 * * * Command’

That would execute the command every three hours. As an asterisk is used for the minutes, the command is executed every minute of these hours.

• ‘<Day>’

The complete sample looks like this:

‘<*|Day[[,Day]|[,Day-Day]. . .]|Day-Day[[,Day]|[,Day-Day]. . .]>’

The value for ‘Day’ has to be choosen in the range from 1 to 31. For a more precise description of the parameter ‘Day’ look at the parameter ‘Minute’ above. Both parameters have the same sample.

Example: ‘* * 5-10,20-25 * * Command’

That would start the command at the fifth to 10. of a month and at the 20. to 25. of a month. As asterisks are used for the other parameters, the command will be executed every minute at these days!

• ‘<Month>’

The complete sample looks like this:

‘<*|Mon[[,Mon]|[,Mon-Mon]. . .]|Mon-Mon[[,Mon]|[,Mon-Mon]. . .]>’

The value for ‘Mon’ has to be choosen in the range from 1 to 12. For a more precise description of the parameter ‘month’ look at the parameter ‘Minute’ above. Both parameters have the same sample.

Example: ‘* * * 1,5-7 * Command’

This would start the command in the months from January to May and in July. It will be executed every minute in these months!

• ‘<Day_of_week>’

The complete sample looks like this:

‘<*|DoW[[,DoW]|[,DoW-DoW]. . .]|DoW-DoW[[,DoW]|[,DoW-DoW]. . .]>’

The value for ‘DoW’ has to be choosen in the range from 1 to 7. 1 corresponds to Monday, 7 corresponds to Sunday. For a more precise description of the parameter ‘Day_of_Week’ look at the parameter ‘Minute’ above. Both parameters have the same sample.

Example: ‘* * * * 1-5 Command’

This would start the command from Monday to Friday every minute, your computer is turned on.

• ‘<Path/<Command|Script>>’

Here you have to enter the command or script that shall be executed. You always have to use the complete path. In the above example in line seven the command ‘newshell’ needs no path, because this is an internal command, which is not loaded from a data media. System commands, which are located in the ‘C:’ drawer likewise don’t need a path. All other commands do.

• examples:

     1. ‘* * * * * "dir >> t:test sys:tools"’

This command is executed every minute.

     2. ‘0,30 8-16 * * 1-5 s:Gong’

This script is executed Mondays to Fridays from eight a.m. to four o’clock p.m. every full hour and every half hour.

     3. ‘0 15 24 12 * execute "s:12. 24."’

This script is execute at three o’clock p.m. the 24.th of December.

     4. ‘0 20 15,16 4 7 s:Example’

This script is executed the 15.th and 16.th of April at eight o’clock p.m. (when this date is a Sunday). Of course two successive days both can’t be Sundays, but this example shall illustrate the AND- combination.

When you want to start scripts, the ‘S- flag’ of the files has to be set (to do this with the script ‘News’, which is in the directory ‘S:’, type the following in a ‘CLI’: ‘protect s:News s add’). The second way to start a script is to use the command ‘execute’ like in the third example (you don’t have to set the ‘S- flag’ here). You have to start commands inside scripts with their full path! You don’t need the path, when you use intern commands or commands in the ‘C:’-directory. When you have problems with commands, not doing what you want them to do, check the commands for spaces. Try quoting the part of the command including spaces. Test the command you wish to use in a ‘Shell’, before you use it in a script.

2.1.3.8 The string gadget ‘Timerinterval’

Here you can select a file that is executed in regular time intervals. This entry therefore can be used to save the project you are working on at the moment every ten minutes. The moments when this happens are always relative to the boot time, i.e. (boot time + n * interval). To change the interval see Section 2.1.4.6 [The sliders ‘Hours’/ ‘Minutes’], page 20.

When you press the gadget in the right of the string gadget, a file requester pops up, with which you can select the desired file. You only can change the entry in this string gadget, when you have selected ‘interval’ or ‘fixed & interval’ at the gadget ‘Type’ on the page ‘Cron Setup’.

Otherwise the gadget will be ghosted. You can select an executeable or a script in ASCII format. When you want to use a script the ‘S- flag’ has to be set. To set this flag for the script ‘Example’ in the directory ‘s:’ type in a ‘CLI’ window: ‘protect s:Example s add’. You have to start commands inside scripts with their full path! You don’t need the path, when you use intern commands or commands in the ‘C:’-directory.

2.1.4 The page ‘CronSetup’

This page is shown, when you make one of the following actions on the graphical user interface:

• Activate the gadget ‘Cron Setup’

• Select the menu ‘Pages⇒Cron Setup’

• Press the keys AMIGA 3

• Press the keys RALT SPACE, when you are on the page ‘Path Setup’

• Press the keys LALT SPACE, when you are on the page ‘Global Statistics’

On this page you can modify the time values, that TGCron accesses. Furthermore you can select the date format, the language, and which type of alarm you want to use. You see three different gadgets and four sliders. You can change the shown values in different ways:

• Slider

− Click with the mouse on a slider and move the mouse, pressing the mouse button, to the left or to the right, the knob follows the mouse. When you reached the desired value, release the mouse button.

− Press the underlined character on your keyboard: The knob moves to the right.

− Press the underlined character on your keyboard together with a SHIFT- key: The knob

moves to the left.

• Gadgets

− Activate the gadget with a mouseclick and hold the mouse button pressed: Select the

desired entry (activation in the text field). − Activate the gadget with a mouseclick: the next entry is selected (activation in the image

field or in the text field) − Press the underlined character on your keyboard: The next entry is selected. − Press the underlined character together with a SHIFT- key: The previous entry is selected.

When you start TimeGuardian the first time, you see the default values. Otherwise you see the last saved configuration.

2.1.4.1 The slider ‘Startuptime’

With this slider you have to enter the time in minutes, your computer needs to boot, rounded to the next full minute. When your system needs for example two minutes and 30 seconds to start, you would have to enter three here. This is necessary, so that the background program can decide whether the computer is switched on or had a reset. The program decides, that the computer was switched on, when the difference between the last saved value and the actual time is greater than the value that you entered with the slider. Otherwise a reset is supposed.

2.1.4.2 The slider ‘Increment’

This value is the time in minutes, when the background program saves, that the computer was still working. The smaller you make this value, the more accurate are the values in the three saved files which you can select on the page ‘Path Setup’. When you enter two, for example, every two minutes is saved, that the computer still is working. The greatest error which may occur at this value is then two. The greater the entered value, the inaccurate is the supposed switch off moment, as the last saved value for a running system is taken.

2.1.4.3 The gadget ‘DateFormat’

With this gadget you can enter nine different formats for the date, that is used, when saving the time in the different files. The format, that is used is the one, that is shown. The following formats can be selected:

     `tt-mm-jj', `tt.mm.jj', `tt/mm/jj',
     `mm-tt-jj', `mm.tt.jj', `mm/tt/jj',
     `jj-mm-tt', `jj-mm-tt', `jj-mm-tt'.

Here stands ‘tt’ for the day of the month, ‘mm’ for the short form of the month and ‘jj’ stands for the last two digits of the year. The short form of the month is saved in the selected language.

2.1.4.4 The gadet ‘Locale’

With this gadget you can select a language for the background program TGCron. The config program is always localized in the language that you selected in the language preferences program (certainly you need an approriate file ‘TimeGuardian.catalog’ installed in your drawer ‘Locale:Catalogs/language’). The program runs in the default language english, when this file is missing, or when you use a workbench version < 2.1. In this distribution only the default language english and german are supported.

You can select six different settings:

• localized:

Use this when you have a localized workench and you want the files for the time functions to be written in this language. You need the appropriate file for your language!

• english:

With this and the following selections, the text that is written in the files, appears in the selected language.

• german, french, italian, spanish: time functions: selected language

When you are intrested, in translating the text for the graphical user interface into one of the above languages or into another, please contact one of the adresses at which you can register for the package, see Section 1.5 [How can I register ?], page 3. When you do this (you translate the complete graphical user interface, the error messages and the text for the time functions) you will get a registered version of the package TimeGuardian. Please contact one of us under any circumstances, befor you start to translate!

2.1.4.5 The gadget ‘Type’

With this gadget you can select, which type of background actions shall be executed. The program TGCron executes the commands/ scripts ‘Once a year’ and ‘Once a day’ and writes the protocol files always when its necessary. This happens not depending on the following selectons. You have several possibilitys to choose:

• none:

That means, that neither the script in the string gadget ‘Timer fixed’ nor the command/ script in the string gadget ‘Timer interval’ is executed.

• fixed:

That means, that only the script in the string gadget ‘Timer fixed’ is executed. You can use as many moments and commands in this script as you want. See Section 2.1.3.7 [The gadget ‘Timer fixed’], page 13, for the contents of such a script.

• interval:

Only the command/ script in the string gadget ‘Timer interval’ will be executed in the selected interval as long as the computer is working.

• fixed & interrval: The two above described commands/ scripts will be executed

2.1.4.6 The sliders ‘Hours’/ ‘Minutes’

Here you can set the interval, in which the command/ script ‘Timer interval’ will be executed regularly. You have to set the minutes and hours seperately. When you set the slider ‘hours’ to 0 and the slider ‘Minutes’ to 15, the command/ script will be executed every quarter of an hour. You can change these two values only, when you selected ‘timer & interval’ or ‘interval’ at the gadget ‘Type’, otherwise the two sliders will be ghosted (see Section 2.1.4.5 [The gadget ‘Type’], page 19).

2.1.5 The gadgets ‘Save’/ ‘Use’

When you press the gadget ‘Save’ or the corresponding key on the keyboard, the program saves the current preferences in a way so that they are used immediately and after booting (The preferences are saved in the directory ‘ENVARC:’ and ‘ENV:’ under the name ‘TimeGuardian.config’).

When you press the gadget ‘Use’ or its corresponding key- shortcut, the preferences are saved in the drawer ‘ENV:’ under the same name like above. The configuration will be used immediately, but is only active as long as the computer has no reset. When the computer boots, the configuration is used, that you saved with the gadget ‘Save’. In contrast to the standard preferences programs TimeGuardian is not quit, when you press one of these two gadgets. When you want to leave the program press the ‘Quit’- gadget.

2.1.6 The gadgets ‘Default’/ ‘Restore’

When you press the gadget ‘Default’, some default values will appear. This are the settings, you see when you start the program the first time after the installation.

With the gadget ‘Restore’ you can cancel the changes, that you have made before. This works as long as you did not use one of the two gadgets ‘Save’ or ‘Use’. You will get the settings back, that you have activated with the gadget ‘Use’ the last time.

2.1.7 The page ‘GlobalStatistics’

This page is shown, when you make one of the folowing actions on the graphical user interface:

• Activate the gadget ‘Global Satistics’

• Select the menu ‘Pages⇒Global Statistics’

• Press the keys AMIGA 4

• Press the keys RALT SPACE, when you are on the page ‘Cron Setup’

• Press the keys LALT SPACE, when you are on the page ‘Daily Statistics’

On this page you find a table on which you can see at which days of the actual year the computer was working how long. When you control different projects with different configurations the times and dates shown correspond to the selected configuration. In the lower field you see the total amout of hours and days for the year. In the upper field you see single days with the times for this day. You can move around in the upper field of the page like follows:

• Press the TAB key and use the keys CURSOR UP or CURSOR DOWN: Each time you press one of the keys the bar moves up or down a day.

• Click with the mouse on the desired day.

• Move the proportional gadget on the right side of the field and then click with the mouse on the desired day.

• Click with the mouse on the gadgets next to the proportional gadget and click with the mouse on the dsired day.

When you have highlighted the day, about which you want more information, with one of the above methods, you can activate the page ‘Daily Statistics’ in different ways:

• Enter the key RETURN (the highlighted day is shown)

• Doubleclick with the mouse on the desired day (this works with a day, that isn’t selected yet, too)

2.1.8 The page ‘DailyStatistics’

This page is shown, when you make one of the folowing actions on the graphical user interface:

• Activate the gadget ‘Daily Satistics’

  • Select the menu ‘Pages⇒Daily Statistics’

• Press the keys AMIGA 5

• Press the keys RALT SPACE, when you are on the page ‘Global Statistics’

• Press the keys LALT SPACE, when you are on the page ‘About’

This page is devided into three areas. In the uppermost you can select with a gadget, what you want to see from the selcted day in the middle of the page:

− OnOff: This shows, when the computer was switched on, how many resets it had and when it was switched off. The accuracy of the switch off time depends on the settings of the slider ‘Increment’, see Section 2.1.4.2 [The slider ‘Increment’], page 18, for further information.

− Starts: Here you can see the times, when your computer had a reset.

In the lowest area you see a summary of the selected day. When you cannot see all entrys together, you can move the proportional gadget the same way like on the page ‘Global Statistics’. To change between the two views ‘OnOff’ an ‘Starts’ you can press the SPACE key.

2.1.9 The gadget ‘RemoveCron’

With this gadget you can stop the background program TGCron. You have to answer a requester, whether you want the program to quit or not. When you select ‘Yes’ TGCron is stopped within the next minute. When you select ‘No’ nothing happens, the program continues.

2.1.10 The gadget ‘Quit’

You can leave the program with this gadget. When you have changed the preferences, but didn’t save them a requester pops up. When you select ‘Yes’ the program is left without saving. When you select ‘No’ or press the RETURN key you’re back in the program. You can save the preferences manually and leave the program afterwards.

2.1.11 The online help

The online help is a feature to get easily and fast help when you use the graphical user interface. To use this feature the ‘amigaguide.library’ has to be installed correctly on your system (located in the drawer ‘Libs:’). When this is not true you can’t use this feature. Where you can obtain the library, see Section 3.3 [Where do I get the Amiga-Guide?], page 31.

When the above described file is installed on your computer, you can get help about an element of the graphical user interface by moving the mouse pointer over the corresponding element (do not click it) and pressing the HELP key afterwards. The corresponding text for the element from this documentation will appear.

2.1.12 The AREXX-Port

At the present version the program only supports the standard MUI commands:

• QUIT

• HIDE

• SHOW

• INFO

• HELP

For more information see MUI.guide from the distribution of MUI.

2.1.13 What do the errormessages mean ?

  • ‘Cannot quit now, still
    some asl popups opened.’

This message appears, when you have to close a file requester, which was opened before, before you can leave the program

• ‘You changed the configuration without saving
Do you really want to quit?’

That remembers you, that you haven’t pressed the gadget ‘Save’ yet. Do this before you leave the program.

• ‘TGCron is not running!’

The background program TGCron is not running yet, therefore the change of the configuration will have no effect! When you want to use the new configuration, you have to start TGCron.

• ‘File already exists!
Do you want to replace it?’

A file you want to save, is already saved with this name. When you select ‘Yes’ the old file is overwritten. When you select ‘No’ you can save the file again and change the name.

• ‘Do you want to remove TimeGuardianCron?’

When you want to stop the background program TGCron select ‘Yes’, otherwise ‘No’.

• ‘You have to define a Log File,
an empty string is not allowed!’

You have not entered anything in a string gadget, this is not allowed her. You have to enter a name for the file.

• ‘Invalid config-file: *’

The file * is not a valid config file, please select an other file.

• ‘Could not load config-file: *
I will use defaults instead!’

The file * could not be opened correctly, while starting TimeGuardian. The program will use the values, you get, when you press the gadget ‘Default’. Change the configuration to suit your needs and press ‘Save’ to correct the error.

2.2 Das Programm TGCron

After setting up your preferences with TimeGuardian, you may start TGCron. According to your setup the program executes the appropriate programs and scripts and stores the runtime information of your computer. Its CPU usage is very low. While waiting for the next event it uses no CPU power, only when beeing started and each minute during a short moment and when it looks for an event or writes some information to disk.

2.2.1 What does TGCron do?

TGCron is designed to stay in memory all the time and run as a background task. It executes user defined events at predefined moments. This means on the one hand that script files are executed, on the other it stores by predefined precission information about the runtime and when your computer has been booted. It is very important to configure TGCron for your personal environment, therewith it is able to accomplish your events. How you can do that, see Section 2.1 [The program

TimeGuardian], page 4.

For further information on the files, refer to Section 2.1.3 [The page ‘Path Setup’], page 7 and Section 2.1.4 [The page ‘Cron Setlup’], page 17.

TGCron searches for the following files during startup:

• These files at every boot:

− ‘ENV:TimeGuardian.config’

TimeGuardian saves the current preferences to this file, TGCron reads them from it. − ‘s:TimeGuardian.key’

This is your personal keyfile. You will get it when you become a registered user. If this file does not exist or has been modified, you will only be able to use TGCron about 1^{1}/_{2}h a day. How to get your key, please refer to Section 1.5 [How can I register for TimeGuardian?], page 3.

− ‘[Log path]TG private/LastBoot’ The time of your last boot is stored here. Don’t change anything here, when you want, that the values are correct in ‘[OnOff]’ and ‘[UpTime]’!

− ‘[Log path]TG private/UpTime’ Here is stored the time in minutes, how long your computer has been running on the current day. Don’t change anything here, when you want, that the values are correct in

‘[UpTime]’!

− ‘[Log path]TG private/UpDate’ This file contains the point of time, when your computer was turned off for the last time. Don’t change anything here, when you want, that the values are correct in ‘[UpTime]’ and

‘[OnOff]’!

− ‘[Log path]TG private/Resets’ Here is the number of resets stored since your last startup. Don’t change anything here, when you want, that the values are correct in ‘[OnOff]’!

− ‘[Log path][LogFile].<yy>’

At every startup or reset the curent time is appended to this file.

• These files only at the first boot of a day or at midnight:

− ‘[Once a day]’

This file will be executed.

− ‘[Log path][UpTime].<yy>’

Once a day the full runtime of the previous day is stored here.

• This file at a cold boot (not at a reset):

‘[Log path][OnOff].<yy>’

The information when the computer has been turned on and how long is stored here during startup.

• This file only at the first boot of the year:

‘[Once a year]’

This file will be executed.

• These files only if selected:

 − ‘[Timer fixed]’

If you selected this at the page ‘Cron Setup’, this file will be read during startup and stored internally. Then it will be checked every minute for a user event.

 − ‘[Timer interval]’

If you selected this at the page ‘Cron Setup’, this file will be executed at fixed intervals relative to the boottime.

Neither change any files in the ‘.../TG_private’ directory nor with a ‘.key’ or ‘.config’ suffix.

The marks have the following meaning:

− If a filename is marked like this, the filename or parts of the filename cannot be influenced by

the user. − A filename marked like ‘[this]’, shows that the user is able to configure the path/filename or

path and filename at the ‘Path Setup’ page. The string gadget in the configuartion program has the same name.

− A mark like this ‘<yy>’ means, that a suffix is appended automatically by program, generated from the last two digits of the current year.

2.2.2 How can I get information about TGCron

If you want to get further information about TGCron, please enter the following from the CLI (replace ‘<path>’ by the directory where you have installed the package to): ‘<Pfad>/TGCron’ There are two different possible messages, according to the fact you have TGCron already started or not.

1. TGCron has already been started:

‘TGCron v1.42 (4.7.1994) by G.Körner/J.Matern
use -r option to remove...’

2. TGCron is not running:

‘TGCron v1.42 (4.7.1994) by G.Körner/J.Matern
use -i option to install...’

You can also get information about the revision and the compilation date by using the ‘version’- command. Enter the following commandline: ‘version full <path>/TGCron’ You should get this result: ‘TGCron v1.42 (4.7.1994)’. If you don’t get that, you are definitely reading the wrong documentation.

2.2.3 How can I start TGCron?

If you didn’t choose the automtic startup option during installation, you have two different options to start TGCron by hand:

1. Start from CLI To start it from CLI please enter the following: ‘<path>/TGCron -i’ As TGCron does not detach from your CLI, you have to use the ‘run’ command to get your shell back after startup as follows: ‘run < nil: > nil: <path>/TGCron -i’. Replace <path> by the directory you have the TimeGuardian package installed to.

2. Workbench startup If you want to start TGCron from your Workbench, just doubleclick its icon.

When you try to start the program, but it is already running, you will get the following results:

− At the CLI:

You will get the message: ‘TGCron is already running’ This shows that TGCron is already running and you cannot restart it again.

− At the workbench

You will get a requester asking you: ‘Do you want to remove TimeGuardianCron?’

If you want to remove it, click on ‘Yes’ or press RETURN. Choosing ‘No’ or pressing ESCAPE does not affect TGCron and it keeps on running. Regardless, what you choose, it won’t be started again!

2.2.4 How can I stop TGCron?

If TGCron is already running, you have three options to remove it:

1. From CLI:

Please enter the following command: ‘<path>/TGCron -r’ Replace <path> by the directory you have the TimeGuardian package installed to. Then you

will get the message: ‘TGCron will be removed after next time interval’. The end of the

time interval is normally the next full minute.

2. At the Workbench:

Just doubleclick its icon. You will get a requester asking you: ‘Do you want to remove

TimeGuardianCron?’

If you want to remove it, click on ‘Yes’ or press RETURN. Choosing ‘No’ or pressing ESCAPE does not affect TGCron and it keeps on running. You won’t get any further messages, but as similar as removing it from CLI it will quit after the next full minute.

3. Start TimeGuardian and press the ‘Remove Cron’ button, refer to Section 2.1.9 [The gadget ‘Remove Cron’], page 22 for detailed information.

2.2.5 What do the error messages mean?

You can get the following error messages from TGCron:

• ‘Invalid .config file. Use TimeGuardian to create a new one!’

TGCron reads data from the file ‘env:TimeGuardian.config’, but it was changed or corrupted and is not valid any more. Start TimeGuardian, enter a new configuration and store it with the ‘Save’ button.

• ‘Could not open env:TimeGuardian.config!
Use TimeGuardian to create it’

Start TimeGuardian, enter your configuration and store it with the ‘Save’ button.

• ‘Could not open locale, please check configuration!
Using english instead’

You selected ‘localised’ at the ‘localisation’ switch, but the program is not able to find

the corresponding files in ‘Locale:languages/your_country/TimeGuardian.catalog’ (re-

place ‘your_country’ by the name of the locale your computer is running with), therefore it will use english as a default. To get rid of this message, install the needed .catalog file or select one of the hard coded languages, ‘english’ for example.

• ‘Your evaluation time has expired for today,
Please consider registering !!’

You have reached the end of your evaluation time for TGCron as an unregistered user. You cannot restart it again on this day. If you want to use the package unlimited, please become a registered user, look at Section 1.5 [How can I register for TimeGuardian?], page 3 for detailed information.

• ‘CronParser ERROR:
Wrong number of Arguments in CronTable line:
CronTab_line
Please correct it!’

The displayed ‘CronTab_line’ from the file that you have choosen at the string gadget ‘Timer fixed’, contains too much or not enough entries. Correct the line to get rid of this requester, because it will be shown again and again. For detailed information on the CronTab, refer to Section 2.1.3.7 [The string gadget ‘Timer fixed’], page 13.

• ‘CronParser ERROR:
Bad entry '*' in CronTable line:
CronTab_line
Please correct it!’

The entry ’*’ in the ‘CronTab_line’ is not valid. Correct the line to get rid of this requester, because it will be shown again and again. For detailed information on the CronTab, refer to Section 2.1.3.7 [The string gadget ‘Timer fixed’], page 13.

• ‘CronParser ERROR:
Value * out of range (*...*)
Please search for commandline and fix it !’

The number after ‘Value’ is not in the range, that is shown in paranthesis. For detailed information on the CronTab, refer to Section 2.1.3.7 [The string gadget ‘Timer fixed’], page 13.

• ‘CronParser ERROR:
Range *-* out of range (*...*)
Please search for commandline and fix it !’

The range after ‘Range’ has to be in the range that is shown in parenthesis. For detailed information on the CronTab, refer to Section 2.1.3.7 [The string gadget ‘Timer fixed’], page 13.

3 Where do I get the support programs?

3.1 Where can I obtain MUI?

You can get MUI directly from the author, where you get a registered version, and can use all the features of the MUI graphical user interface (see Section 5.1 [Where can I register for MUI?], page 36). Or you obtain the Fish-Disks 967/968 from your local public domain dealer or the

Aminet.

When you think, it is not worth to install MUI, because of one single program, this may be correct. But the background program TGCron doesn’t use MUI at all. Only the preferences program uses it for its graphical user interface. So MUI is not started all the time but only when you want to change the configuration. So you can install this package even on computers with little menmory, because you don’t need MUI all the time.

When you have not installed MUI yet, you have to do this before the installation of TimeGuardian by all means, as the installation will be aborted otherwise. See Chapter 4 [How can I install TimeGuardian?], page 32, for further information.

3.2 Where can I obtain Installer?

For the easy installation of the program package, you need the program Installer from C*mm*d*re. When you have this program not installed on your computer, you can obtain it on different ways.

• With the Workbench 2.0/2.1/3.0

  • On the Fish-Disk 870

• On the distribution of MUI.

• On most distributions of commercial software

3.3 Where can I obtain Amiga-Guide?

When you want to use the online help, you need the ‘amigaguide.library’ in your ‘LIBS:’- drawer (see Section 2.1.11 [The online help], page 23). You don’t need it to run the package TimeGuardian, but won’t be able to use the online help feature.

The ‘amigaguide.library’ is on the Fish-Disk 870 and is distributed with Workbench version 3.0.

4 How can I install TimeGuardian?

This chapter describes the installation of the package TimeGuardian. When you want to install other packages, not included in this distribution (see Section 1.4 [What do I need to use TimeGuardian], page 2), look in the corresponding distributions for installing them.

You have two possibilities to install the package TimeGuardian. You should use the first one:

1. You use the Installer script, that is in this package (the program Installer has to be in the search path e.g. ‘C:’). For a precise description of the script , see Section 4.1 [The installation with Installer], page 32. When you have not installed the Installer, you can obtain it, see Section 3.2 [Where do I get the Installer?], page 30.

2. You can install the package manually, for a precise description how do do this, see Section 4.2 [The installation withoutInstaller], page 34.

4.1 The installation with Installer

You have two choices to start the installation script:

1. Doubleclick with the mouse the icon ‘TG.install.deutsch’:

The installation starts. The text appears german. When you want to use the geramn text please read the german documentation.

2. Doubleclick with the mouse the icon ‘TG.install.english’:

The installation starts. The text appears in english.

Now follows a step by step description of the installation:

• At the beginning you have the possibility to choose the user level for the installation. When you select ‘Novice’ you won’t be able to select anything more. You cannot control, where the package is installed. All files which are necessary to work corretly will be copied, and the preferences program will be started. At the end you get a summary about what was saved to which location. To quit the installation you have to leave the preferences program. When you want to install the package press ‘Proceed With Install’. With ‘Help...’ you get some information.

• When you select the user level ‘EXPERT’ you can choose whether you want to install the package (‘Install for Real’) or whether you just want to test the installation script (‘Pretend to

Install’).

Furthermore you can choose, whether all actions are logged or not. You can write the file ‘Sys:TGInstall.e.log’ to disk (‘Log File’), you can print the results (‘Printer’) or you can suppress a log file (‘None’). The installation is continued with ‘Proceed’.

• The installation checks, whether MUI is installed with the correct version. When this is not true it is aborted.

• Now you can choose what you want to install.

The first gadget means, that only the necessary programs and config files are copied. The second gadget means the documentation. Here are three files at the moment: − The file ‘TimeGuardian.dvi’, that you can print or view, when you have installed TEX. − The file ‘TimeGuardian.guide’, that you can read with AmigaGuide. This file con-

tains the information for the online help. This file is always copied, when the file ‘Libs:amigaguide.library’ is installed on your system.

− The file ‘TimeGuardian.asc’, you can read with any text viewer or editor, or that you

can print.

• When you have activated the first gadget:

− You are asked, if the background program TGCron should be started every time the com-

puter is booting. − When you answered the last question with ‘Yes’, the program checks whether this is

possible on your system. − When multiple possibilities exist, to start the background program, you can select which of

them you prefer. When you select ‘From CLI (user-startup)’ the file ‘S:user-startup’ will be changed in an appropriate way. When you select ‘From WB (WBStartup)’ the icon of the progam TGCron is copied to the drawer ‘Sys:WBStartup’. Users, that work mainly on the workbench should use the second possibility.

− Now you can choose, where the programs shall be copied to. The installation makes a proposal. When you want a new drawer for the package created, select ‘Make New Drawer...’, otherwise choose an existing one.

− You can continue with the button ‘Proceed’. − Even if you did not select the documents to be installed the file

‘TimeGuardian.guide’ will be copied, when the file ‘Libs:amigaguide.library’ is in-

stalled. When this is not true a message appears, that you won’t be able to use the online help.

• When you have selected the second gadget:

− You now have to choose a drawer where the documents shall be copied to. The program

suggests a subdirectory, in the drawer, where you copied the programs to.

 − The file ‘TimeGuardian.guide’ is only copied, when the file ‘Libs:amigaguide.library’

is installed on your system. When this is not true a message appears, that you won’t be able to use the online help.

• When you have installed the programs you now can decide, whether you wish to start the preferences program TimeGuardian. Start the program with selecting ‘Yes’. Change the configuration to suit your needs and press the gadget ‘Save’ before you leave the program with the gadget ‘Quit’. When you don’t start the program, you have to answer a message, that you will have to start the preferences program later on. Select ‘Proceed’ here to leave the installation correctly.

• When you have left the preferences program or you answered the message, the installation is complete and the last message appears.

4.2 The installation without Installer

To install the package without the Installer please follow these steps:

1. Open a Shell window (doubleclick the icon ‘Shell’ in the drawer ‘System’ of your boot partition).

2. Please enter ‘version libs:muimaster.library’. You should get this result:

‘libs:muimaster.library 7.973’. When you get a smaller number or ‘object not found’,

you have to install MUI version 2.0 or greater before you can continue.

3. When you want to install the package in a drawer, that already exists, please enter ‘cd drawer’ (instead of drawer please use the name of the directory). When you want to copy the files to a new drawer, then you have to create it with the command ‘makedir drawer’, then change to this drawer with ‘cd drawer’ (please use the name of the new directory instead of drawer in the two command lines, e.g. TimeGuardian).

4. Now put the disk with the package into any drive (when you have copied the files already to your hard disk then enter: ‘assign TimeGuardian: sourcedrawer’, where sourcedrawer is the drawer, where the files are on your hard disk).

5. To copy the programs enter: ‘copy TimeGuardian:bin/#? ""’. Some messages appear, which files have been copied already.

6. When you want to copy the locale files (only german available at the moment) please enter:

‘copy TimeGuardian:Catalogs/deutsch/#? Locale:Catalogs/deutsch’. You should get a

message about a copied file.

7. When you want to copy the english documentation files please enter

‘copy TimeGuardian:doc/english/#? Help:english’.

8. Please enter ‘copy TimeGuardian:env/#? envarc:’ and ‘copy TimeGuardian:env/#? env:’

now.

9. Please enter ‘copy TimeGuardian:s/#? s:’.

10. When the background program shall be started, while booting, please enter:

− start from Workbench:

‘copy TimeGuardian:bin/TGCron.info Sys:WBStartup’

Open the drawer ‘WBStartup’ on the Workbench and click at the icon ‘TGCron’. Activate the menu ‘Icons⇒Information...’ or press the keys AMIGA I. A window should pop up. Now enter in the string gadget ‘Default Tool:’: ‘path/TGCron’ (you have to replace path with the full path, where you copied the programs to). Leave the window with the gadget ‘Save’. or:

− start from user-startup:

Add the following lines to the file ‘s:startup-sequence’ or ‘s:user-startup’: ‘run <

nil: >nil: path/TGCron’ (you have to replace path with the full path, where you copied the programs to). You can do this with the help of an editor.

11. When you want to start the preferences program from the same drawer as the standard prefe-

rences programs then enter: ‘copy TimeGuardian:bin/TimeGuardian.info Sys:Prefs’. You

have to change the ‘Default Tool:’ like described above at step 10 (from Workbench). Enter

the name ‘path/TimeGuardian’ instead of ‘path/TGCron’!

12. To use your personal configuration start the program TimeGuardian and leave it with the gadget ‘Save’.

It may be, that you run into problems with this above described installation. This can be, when the assign ‘Help’ and the drawer ‘Help:english’ don’t exist. You can copy the documentation files to any other drawer you want. You have to change the file ‘Envarc:AmigaGuide/path’ and ‘Env:AmigaGuide/path’ so, that the drawer, where you copied the file ‘TimeGuardian.guide’ to, is searched. E.g. add the line ‘dh0:docs’ when you copied the above file to the drawer ‘dh0:docs’.

5 What is MUI?

MUI is program package for users and developpers, that makes a graphical user interface available to the programmer and user, with many features. TimeGuardian ist not able to run without MUI. More precise, the preferences program TimeGuardian! Therefore you have installed MUI already or you have to do it before the installation of TimeGuardian. Where you can obtain MUI, see Section 3.1 [Where can I obtain MUI?], page 30.

5.1 Where can I register for MUI?

This application uses MUI - MagicUserInterface \begin{matrix}c \\ ◯ \end{matrix} Copyright 1993 by Stefan Stuntz

MUI is a system to generate and maintain graphical user interfaces. With the aid of a preferences program, the user of an application has the ability to customize the outfit according to his personal taste.

MUI is distributed as shareware. To obtain a complete package containing lots of examples and more information about registration please look for a file called ‘muiXXusr.lha’ (XX means the latest version number) on your local bulletin boards or on public domain disks.

If you want to register directly, feel free to send DM 20.- or US-$ 15.- to Stefan Stuntz Eduard-Spranger-Strasse 7 80935 München GERMANY Index 37

Index

AREXX-Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 assumptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Cron Setup, page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 Daily Statistics, page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Date Format, gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Default, gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 description of the installation . . . . . . . . . . . . . . . . . . . . . . 32 Description, TGCron . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 Error messages, TGCron . . . . . . . . . . . . . . . . . . . . . . . . . . 28 errormessages, TimeGuardian . . . . . . . . . . . . . . . . . . . . . 23 Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 function, package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 gadget Date Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 gadget Default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 gadget Locale. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gadget Quit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 gadget Remove Cron. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 gadget Restore. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 gadget Save. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 gadget Type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gadget Use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Global Statistics, page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Hours, slider . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 Increment, slider . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 Informationen, TGCron . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Installation, with Installer . . . . . . . . . . . . . . . . . . . . . . . . . 32 Installation, without Installer . . . . . . . . . . . . . . . . . . . . . . 34 installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Locale, gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 Menu, Edit=>Last Saved. . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 menu, Project=>Ikonify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 menu, Project=>Open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 menu, Project=>Save as . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 menu, Project=>User . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 menu, Settings=>Create Icons?. . . . . . . . . . . . . . . . . . . . . . 7 menus TimeGuardian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Minutes, slider . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 MUI, about. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

Once a day, string gadget . . . . . . . . . . . . . . . . . . . . . . . . . . 11
Once a year, string gadget . . . . . . . . . . . . . . . . . . . . . . . . . 12
online help. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23
OnOff, string gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10
page About . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
page Cron Setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17
page Daily Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22
page Global Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21
page Path Setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
Program, TimeGuardian . . . . . . . . . . . . . . . . . . . . . . . . . . . 4
Programdescription . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4
Programm, TGCron . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24
programs, other. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30
Quit, TGCron . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22
Quit, TimeGuardian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22
Registrierung . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3
Remove Cron, gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22
Restore, gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20
Runtime, string gadget. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10
Save, gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20
Sharewaregebühr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3
slider Hours . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20
slider Increment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18
slider Minutes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20
slider Startuptime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18
start TimeGuardian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4
Start, TGCron . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27
Starts, string gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9
Startuptime, slider . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18
Stop, TGCron . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28
string gadget Log Path. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8
string gadget Once a day . . . . . . . . . . . . . . . . . . . . . . . . . . 11
string gadget Once a year . . . . . . . . . . . . . . . . . . . . . . . . . . 12
string gadget OnOff. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10
string gadget Runtime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10
string gadget Starts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9
string gadget Timer fixed . . . . . . . . . . . . . . . . . . . . . . . . . . 13
string gadget Timer interval . . . . . . . . . . . . . . . . . . . . . . . 16
Timer fixed, string gadget. . . . . . . . . . . . . . . . . . . . . . . . . . 13

Index 38

Timer interval, string gadget . . . . . . . . . . . . . . . . . . . . . . . 16 Type, gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

Use, gadget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

Appendix A Registration form

------------------------------------------------------------------
REGISTRATIONS-FORM
Firm:              ___________________________________________
First name, Name:  ___________________________________________
Street:            ___________________________________________
Location:          ___________________________________________
Post code:         ___________
Country:           ___________________________________________
Telephone:         _____________
I want to become a registered user of the program package TimeGuardian
and therefore
o sent 20.- DM in notes.
o sent 15.- US-$ in notes.
Please send me the newest version of the above program package and my
personal key file to the above address.
_______________________________________________
   Location, Date                         Signature
------------------------------------------------------------------

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 What is TimeGuardian? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 The features of TimeGuardian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 What can I do with TimeGuardian? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 What do I need, to be able to use TimeGuardian? . . . . . . . . . . . . . . . . . 2 1.5 How can I register for TimeGuardian? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 The programs of the package TimeGuardian . . . . . . . . . 4

2.1 The program TimeGuardian . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2.1.1 The menus of the program TimeGuardian . . . . . . . . . . . . . . . . 5 2.1.1.1 The menu ‘Project⇒Open’ . . . . . . . . . . . . . . . . . . . . . 5 2.1.1.2 The menu ‘Project⇒Save as’ . . . . . . . . . . . . . . . . . . 6 2.1.1.3 The menu ‘Project⇒User’ . . . . . . . . . . . . . . . . . . . . . 6 2.1.1.4 The menu ‘Project⇒Iconify’. . . . . . . . . . . . . . . . . . 7 2.1.1.5 The menu ‘Edit⇒Last Saved’ . . . . . . . . . . . . . . . . . . 7 2.1.1.6 The menu ‘Settings⇒Create Icons?’ . . . . . . . . . . 7

2.1.2 The page ‘About’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.1.3 The page ‘Path Setup’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.1.3.1 The string gadget ‘Log Path’ . . . . . . . . . . . . . . . . . . . . 8 2.1.3.2 The string gadget ‘Starts’ . . . . . . . . . . . . . . . . . . . . . . 9 2.1.3.3 The string gadget ‘Runtime’. . . . . . . . . . . . . . . . . . . . 10 2.1.3.4 The string gadget ‘OnOff’ . . . . . . . . . . . . . . . . . . . . . . 10 2.1.3.5 The string gadget ‘Once a day’ . . . . . . . . . . . . . . . . . 11 2.1.3.6 The string gadget ‘Once a year’. . . . . . . . . . . . . . . . 12 2.1.3.7 The string gadget ‘Timer fixed’ . . . . . . . . . . . . . . . 13 2.1.3.8 The string gadget ‘Timer interval’ . . . . . . . . . . . . 16

2.1.4 The page ‘Cron Setup’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.1.4.1 The slider ‘Startuptime’ . . . . . . . . . . . . . . . . . . . . . . 18 2.1.4.2 The slider ‘Increment’ . . . . . . . . . . . . . . . . . . . . . . . . . 18 2.1.4.3 The gadget ‘Date Format’. . . . . . . . . . . . . . . . . . . . . . 18 2.1.4.4 The gadet ‘Locale’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.1.4.5 The gadget ‘Type’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 2.1.4.6 The sliders ‘Hours’/ ‘Minutes’ . . . . . . . . . . . . . . . . . 20

2.1.5 The gadgets ‘Save’/ ‘Use’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.1.6 The gadgets ‘Default’/ ‘Restore’ . . . . . . . . . . . . . . . . . . . . . . . 20 2.1.7 The page ‘Global Statistics’ . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.1.8 The page ‘Daily Statistics’ . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.1.9 The gadget ‘Remove Cron’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.1.10 The gadget ‘Quit’. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.1.11 The online help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 2.1.12 The AREXX-Port . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 2.1.13 What do the errormessages mean ?. . . . . . . . . . . . . . . . . . . . . 23

2.2 Das Programm TGCron . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 2.2.1 What does TGCron do? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 2.2.2 How can I get information about TGCron . . . . . . . . . . . . . . . . 26 2.2.3 How can I start TGCron? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 2.2.4 How can I stop TGCron?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 2.2.5 What do the error messages mean? . . . . . . . . . . . . . . . . . . . . . . 28

3 Where do I get the support programs? . . . . . . . . . . . 30

3.1 Where can I obtain MUI? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.2 Where can I obtain Installer? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 3.3 Where can I obtain Amiga-Guide? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

4 How can I install TimeGuardian? . . . . . . . . . . . . . . . . . . . . 32

4.1 The installation with Installer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 4.2 The installation without Installer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

5 What is MUI?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

5.1 Where can I register for MUI?. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

Appendix A Registration form . . . . . . . . . . . . . . . . . . . . . . . 39